Binomial probability lookup
- P(X = 3)
- 0.1172
- P(X ≤ 3)
- 0.1719
- P(X < 3)
- 0.0547
- P(X ≥ 3)
- 0.9453
- P(X > 3)
- 0.8281
X ~ Binomial(n = 10, p = 0.5). Exact values, not a normal approximation. Mean np = 5, SD √(np(1−p)) = 1.5811.
| k | P(X = k) | P(X ≤ k) |
|---|---|---|
| 0 | 0.0010 | 0.0010 |
| 1 | 0.0098 | 0.0107 |
| 2 | 0.0439 | 0.0547 |
| 3 | 0.1172 | 0.1719 |
| 4 | 0.2051 | 0.3770 |
| 5 | 0.2461 | 0.6230 |
| 6 | 0.2051 | 0.8281 |
| 7 | 0.1172 | 0.9453 |
| 8 | 0.0439 | 0.9893 |
| 9 | 0.0098 | 0.9990 |
| 10 | 0.0010 | 1.0000 |
Convention and assumptions
X counts successes in n independent trials with a common success probability p. Each trial has two outcomes, and n is fixed. The lookup above computes exact probabilities for any n up to 1,000. The printed tables below give cumulative probabilities P(X ≤ k), rounded to four decimals. A printed 0.0000 or 1.0000 can reflect rounding rather than exact impossibility or certainty.
Sampling without replacement from a small finite population changes the probability across draws; use the hypergeometric reference instead. Shared batch conditions or repeated users may violate independence.
Cumulative probabilities: n = 5
| k | p = 0.05 | p = 0.1 | p = 0.2 | p = 0.3 | p = 0.5 |
|---|---|---|---|---|---|
| 0 | 0.7738 | 0.5905 | 0.3277 | 0.1681 | 0.0312 |
| 1 | 0.9774 | 0.9185 | 0.7373 | 0.5282 | 0.1875 |
| 2 | 0.9988 | 0.9914 | 0.9421 | 0.8369 | 0.5000 |
| 3 | 1.0000 | 0.9995 | 0.9933 | 0.9692 | 0.8125 |
| 4 | 1.0000 | 1.0000 | 0.9997 | 0.9976 | 0.9688 |
| 5 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 1.0000 |
Cumulative probabilities: n = 10
| k | p = 0.05 | p = 0.1 | p = 0.2 | p = 0.3 | p = 0.5 |
|---|---|---|---|---|---|
| 0 | 0.5987 | 0.3487 | 0.1074 | 0.0282 | 0.0010 |
| 1 | 0.9139 | 0.7361 | 0.3758 | 0.1493 | 0.0107 |
| 2 | 0.9885 | 0.9298 | 0.6778 | 0.3828 | 0.0547 |
| 3 | 0.9990 | 0.9872 | 0.8791 | 0.6496 | 0.1719 |
| 4 | 0.9999 | 0.9984 | 0.9672 | 0.8497 | 0.3770 |
| 5 | 1.0000 | 0.9999 | 0.9936 | 0.9527 | 0.6230 |
| 6 | 1.0000 | 1.0000 | 0.9991 | 0.9894 | 0.8281 |
| 7 | 1.0000 | 1.0000 | 0.9999 | 0.9984 | 0.9453 |
| 8 | 1.0000 | 1.0000 | 1.0000 | 0.9999 | 0.9893 |
| 9 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 0.9990 |
| 10 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 1.0000 |
Cumulative probabilities: n = 15
| k | p = 0.05 | p = 0.1 | p = 0.2 | p = 0.3 | p = 0.5 |
|---|---|---|---|---|---|
| 0 | 0.4633 | 0.2059 | 0.0352 | 0.0047 | 0.0000 |
| 1 | 0.8290 | 0.5490 | 0.1671 | 0.0353 | 0.0005 |
| 2 | 0.9638 | 0.8159 | 0.3980 | 0.1268 | 0.0037 |
| 3 | 0.9945 | 0.9444 | 0.6482 | 0.2969 | 0.0176 |
| 4 | 0.9994 | 0.9873 | 0.8358 | 0.5155 | 0.0592 |
| 5 | 0.9999 | 0.9978 | 0.9389 | 0.7216 | 0.1509 |
| 6 | 1.0000 | 0.9997 | 0.9819 | 0.8689 | 0.3036 |
| 7 | 1.0000 | 1.0000 | 0.9958 | 0.9500 | 0.5000 |
| 8 | 1.0000 | 1.0000 | 0.9992 | 0.9848 | 0.6964 |
| 9 | 1.0000 | 1.0000 | 0.9999 | 0.9963 | 0.8491 |
| 10 | 1.0000 | 1.0000 | 1.0000 | 0.9993 | 0.9408 |
| 11 | 1.0000 | 1.0000 | 1.0000 | 0.9999 | 0.9824 |
| 12 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 0.9963 |
| 13 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 0.9995 |
| 14 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 1.0000 |
| 15 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 1.0000 |
Cumulative probabilities: n = 20
| k | p = 0.05 | p = 0.1 | p = 0.2 | p = 0.3 | p = 0.5 |
|---|---|---|---|---|---|
| 0 | 0.3585 | 0.1216 | 0.0115 | 0.0008 | 0.0000 |
| 1 | 0.7358 | 0.3917 | 0.0692 | 0.0076 | 0.0000 |
| 2 | 0.9245 | 0.6769 | 0.2061 | 0.0355 | 0.0002 |
| 3 | 0.9841 | 0.8670 | 0.4114 | 0.1071 | 0.0013 |
| 4 | 0.9974 | 0.9568 | 0.6296 | 0.2375 | 0.0059 |
| 5 | 0.9997 | 0.9887 | 0.8042 | 0.4164 | 0.0207 |
| 6 | 1.0000 | 0.9976 | 0.9133 | 0.6080 | 0.0577 |
| 7 | 1.0000 | 0.9996 | 0.9679 | 0.7723 | 0.1316 |
| 8 | 1.0000 | 0.9999 | 0.9900 | 0.8867 | 0.2517 |
| 9 | 1.0000 | 1.0000 | 0.9974 | 0.9520 | 0.4119 |
| 10 | 1.0000 | 1.0000 | 0.9994 | 0.9829 | 0.5881 |
| 11 | 1.0000 | 1.0000 | 0.9999 | 0.9949 | 0.7483 |
| 12 | 1.0000 | 1.0000 | 1.0000 | 0.9987 | 0.8684 |
| 13 | 1.0000 | 1.0000 | 1.0000 | 0.9997 | 0.9423 |
| 14 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 0.9793 |
| 15 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 0.9941 |
| 16 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 0.9987 |
| 17 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 0.9998 |
| 18 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 1.0000 |
| 19 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 1.0000 |
| 20 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 1.0000 |
Read the requested tail
For whole-number endpoints, P(X ≥ k) = 1 − P(X ≤ k − 1). An exact probability is P(X ≤ k) − P(X ≤ k − 1), and P(a ≤ X ≤ b) = P(X ≤ b) − P(X ≤ a − 1). Treat the CDF below zero as zero. Subtracting rounded cells can lose a final decimal; use full-precision probabilities for final calculations.
Worked example
For n = 10 and p = 0.3, P(X ≤ 3) = 0.6496. Thus P(X ≥ 4) ≈ 0.3504. The exact probability P(X = 3), computed from the mass formula, is 0.2668 rounded. These are model probabilities, not evidence that real trials meet the assumptions.
The expected count is np = 3 and its standard deviation is √[np(1 − p)] ≈ 1.4491. An observed count need not equal its expectation.
Continue
Read the binomial lesson. NIST documents the model and probability functions. The general event-probability calculator does not calculate this distribution.